the sum of two angle measures 130 degrees the difference in the two angle measures is 12 degrees what are two angle measures ?

Found 2 solutions by ikleyn, Boreal:
Answer by ikleyn(52812)

.

a + b = 130,
a - b = 12.

Add the two equations (both sides).


Answer by Boreal(15235)

x and y are the angles
x+y=130
x-y=12
2x=142 by adding
x=71 degrees
y=130-71=59 degrees.